OpenStudy (anonymous):

10 sqr root of 162 - sqr root of 405 + sqr root of 200 - 10 sqr root of 500

OpenStudy (lilg132):

\[224 - 90\sqrt{2} - 9\sqrt{5} + 10\sqrt{2} = -103\]

OpenStudy (anonymous):

simplify not solve, sorry

OpenStudy (lilg132):

\[90\sqrt{2} - 9\sqrt{5} + 10\sqrt{2} - 10\sqrt{500}\]